x+5y=4
x=5y-16
solve equation by the substitution method

Answer by Ave(106)

insert all of one equation into another
x=5y-16, you already know what x is in this equation so plug that into the other one.
(5y-16)+5y=4 combine like terms
10y-16=4 add 16 to both sides
10y=20 divide by 10
y=2
Now that you know y is 2, find x by plugging in 2 for y in either equation
x=5(2)-16
x=10-16
x=-6
(-6,2)